### v4

# PaDeLLM-NER: Parallel Decoding in Large Language Models for Named Entity Recognition

Abstract
In this study, we aim to reduce generation latency for Named Entity Recognition (NER) with Large Language Models (LLMs). The main cause of high latency in LLMs is the sequential decoding process, which autoregressively generates all labels and mentions for NER, significantly increase the sequence length. To this end, we introduce Parallel Decoding in LLM for NER (PaDeLLM-NER), a approach that integrates seamlessly into existing generative model frameworks without necessitating additional modules or architectural modifications. PaDeLLM-NER allows for the simultaneous decoding of all mentions, thereby reducing generation latency. Experiments reveal that PaDeLLM-NER significantly increases inference speed that is 1.76 to 10.22 times faster than the autoregressive approach for both English and Chinese. Simultaneously it maintains the quality of predictions as evidenced by the performance that is on par with the state-of-the-art across various datasets.
[…]
PaDeLLM-NER: Parallel Decoding in Large Language Models for Named Entity Recognition
**footnotetext: Corresponding author.
1 Introduction
Variant Input Unstructured Text Output Structured Label-mention String
Augmented Language Paolini et al. (2020); Das et al. (2023)
Japan, co-hosts of the World Cup in 2002 and ranked 20th in the world by FIFA, are favourites to regain their title here.
[Japan — LOC], co-hosts of the [World Cup — MISC] in 2002 and ranked 20th in the world by [FIFA — ORG], are favourites to regain their title here.
Structured Annotation Lu et al. (2022, 2023a); Wang et al. (2023c)
Cuttitta announced his retirement after the 1995 World Cup , where he took issue with being dropped from the Italy side that faced England in the pool stages.
((PER): (Cuttitta), (MISC): (1995 World Cup), (LOC): (Italy), (LOC): (England), (ORG): (NULL))
Table 1: Structured output string format used in the literature. The examples come from CoNLL2003 dataset.
Named Entity Recognition (NER), a fundamental task in Natural Language Processing (NLP), aims to extract structured information from unstructured text data. This includes identifying and categorizing key elements such as Organization, Geopolitical Entity and so on (referred to as “labels”) in inputs, and pairing them with relevant text spans extracted from the text (termed “mentions”). Conventionally, NER tasks are carried out through an extractive paradigm that entails token-level classification and the subsequent extraction of identified tokens Ma et al. (2020); Liu et al. (2021).
Recent advancements in Large Language Models (LLMs) Raffel et al. (2020a); Muennighoff et al. (2022); Touvron et al. (2023a, b); Bai et al. (2023); Gao et al. (2023); Liu et al. (2023d); Yang et al. (2023a) have revolutionized numerous foundational tasks in NLP, including NER tasks Paolini et al. (2020); Lu et al. (2022); Das et al. (2023); Lu et al. (2023a); Wang et al. (2023c), through the adoption of a generative paradigm. This paradigm involves instruction-tuning a sequence-to-sequence (seq2seq) model. The model takes a sequence of unstructured text as input and produces a sequence of structured label-mention pairs as output. Generally, the output structured string should be formatted to meet two criteria: (1) it should have a clear and straightforward structure that facilitates post-processing for label and mention extraction, and (2) it needs to be generated fluidly and efficiently from the perspective of language models Wang et al. (2023b).
In Table 1, we list two typically used autoregressive output formats found in the literature : (1) accommodate original input text to contain label information, which is referred to as “augmented language” Paolini et al. (2020); Das et al. (2023); (2) directly using a customized, easily-parsed structured format to output all labels and mentions, which is called “structured annotation” Lu et al. (2022, 2023a); Wang et al. (2023c). These formats present certain challenges. For example, augmented language necessitates duplicating all original input text, thereby increasing output length and resulting in inference inefficiency. While structure annotation avoids replicating the entire input, it produces all labels and mentions in an autoregressive manner. This implies that each subsequently generated pair depends on its preceding pairs, and when the number of label-mention pairs is large, it will lead to longer sequences. As demonstrated in Chen et al. (2023c); Ning et al. (2023), high latency in LLMs mainly stems from lengthy sequence generation, we believe that by reducing the length of sequence, a more efficient inference scheme can be provided for NER tasks.
In light of this, we propose Parallel Decoding in LLM for NER (PaDeLLM-NER), a novel approach to accelerate the inference of NER tasks for LLMs. PaDeLLM-NER empowers the model with the capability to predict a single label-mention pair within a single sequence, subsequently aggregating all sequences to generate the final NER outcome. Specifically, in the training phase, we reconstruct the instruction tuning tasks, enabling LLMs to predict the count of mentions for a specific label and to identify the $n^{th}$ mention within the entire input for that label (Figure 1). In the inference phase, LLMs first predict the number of mentions for all labels, then predict all label-mention pairs in parallel (Figure 2). Finally, results from all sequences are aggregated and duplicate mentions across labels are eliminated based on prediction probability. This approach results in a more efficient inference method for NER tasks, producing shorter output strings and enabling parallel decoding label-mention pairs in batches.
Comprehensive experiments have been conducted, demonstrating that PaDeLLM-NER effectively reduces the number of tokens produced in each sequence, thereby decreasing inference latency. Additionally, it maintains or even enhances prediction quality in both flat and nested NER for English and Chinese languages, compared to existing methods in the literature. To conclude, our contributions are as follows:
  • We present PaDeLLM-NER, a novel approach tailored for NER using LLMs. This approach can predict all label-mention pairs in parallel, effectively reducing inference latency.
  • Extensive experiments have been conducted, revealing that PaDeLLM-NER significantly improves inference efficiency. By completely decoupling the generation of label-mention pairs, the average sequence length is reduced to around 13% of that produced by conventional autoregressive methods. Correspondingly, the inference speed is 1.76 to 10.22 times faster than these previous approaches.
  • Comprehensive experiments demonstrate that, in addition to its enhanced prediction speed, PaDeLLM-NER also maintains or surpasses the prediction quality of conventional autoregressive methods, on par with state-of-the-art performance on many NER datasets.
To the best of our knowledge, our techinique stands as a pioneering approach in accelerating NER inference in LLMs by parallel decoding all label-mention pairs. This unique characteristic makes it complementary to other inference acceleration methods such as LLM.int8() Dettmers et al. (2022) and speculative sampling Chen et al. (2023a); Leviathan et al. (2023). Thus, it can be efficiently integrated with these methods.***Code is available at URL masked for anonymous review.
2 Related Work
2.1 Generative Models for NER
Before the era of LLMs, most research approached NER as a sequence labeling task, where each token is assigned a pre-defined tag (e.g., BIO scheme). In this line of work, usually pre-trained transformer-based language models Ma et al. (2020); Liu et al. (2021) is combined with a tailored prediction head to perform a token-level classification, followed by the extraction of identified tokens.
Encouraged by the success of unifying multiple NLP tasks into a single seq2seq paradigm Brown et al. (2020); Lester et al. (2021), especially with the evolution of LLMs Raffel et al. (2020b); Achiam et al. (2023); Touvron et al. (2023a); Yang et al. (2023a), the trend of applying seq2seq models to NER tasks is gaining momentum Xu et al. (2023), with both inputs and outputs being represented as sequences of text Paolini et al. (2020); Lu et al. (2022); Das et al. (2023); Lu et al. (2023a); Wang et al. (2023c). Recently, the focus of work on NER using LLMs has shifted towards zero-shot Xie et al. (2023); Sainz et al. (2023) or few-shot learning Ashok and Lipton (2023); Chen et al. (2023b); Das et al. (2023); Wang et al. (2023b), utilizing in-context learning Chen et al. (2023b); Wang et al. (2023b), self-consistency Wang et al. (2022b); Xie et al. (2023) or learning programming Friedman et al. (2023); Sainz et al. (2023).
Unlike previous studies emphasizing few-shot performance with training-free prompt learning Liu et al. (2023c); Lu et al. (2023b), our work focus on a fully supervised setting. More importantly, our primary objective is to speed up NER inference.
2.2 Inference Speedup in LLMs
Modern LLMs employ a sequential decoding strategy for token generation, which poses a significant challenge in terms of parallelization, especially as model size and sequence length increase Ning et al. (2023). There is plenty of work in the literature to address this challenge Wang et al. (2021); Frantar et al. (2023); Santilli et al. (2023); Xiao et al. (2023). One line of work falls into training-free category such as introducing extra modules for speculative sampling Chen et al. (2023a); Leviathan et al. (2023). Another approaches explore modifying model architecture to accelerate inference, such as exiting at earlier layer Elbayad et al. (2019); Schuster et al. (2022), or designing entirely different training and inference mechanisms Lan et al. (2023); Yang et al. (2023b); Zhang et al. (2023). Different from previous works, we focus on exploring the inference speedup in LLMs with a focus on the NER task without the change of model architecture or modules.
3 Method
Figure 1: PaDeLLM-NER training paradigm: texts within frames of the same color represents one training example, where texts inside the solid-line frame are the input, and those inside the dashed-line frame are the output. Italic texts are prompt templates. The “entity type” signifies the label being predicted. The “¡num¿” indicates count of mentions for that label, and “¡mention n¿” refers to the $n^{th}$ mention of a label in the input.
Figure 2: PaDeLLM-NER inference paradigm: texts enclosed in frames with identical colors indicate sequences of the same label. Specifically, the texts within solid-lined frames represent the added templates, while those within dashed-lined frames denote the prediction. In Step 1, the model predicts the number of mentions for all labels while in Step 2, it predicts the mentions. By aggregating mentions and labels from all sequences, the final NER results are obtained. Duplicate mentions appearing in different labels are resolved using prediction probabilities.
In this section, we delve into the details of PaDeLLM-NER. First, we focus on reframing the instruction tuning tasks as outlined in Section 3.1. Second, we explore the two-step inference process, detailed in Section 3.2. Finally, we discuss the aggregation of results and the technique for eliminating duplicate mentions across labels, which is elaborated in Section 3.3. An illustration of PaDeLLM-NER is shown in Figure 1 and Figure 2.
3.1 Reframing of Instruction Tuning
Illustration of the reframing is presented in Figure 1. As an example, we use a case from the CoNLL2003 dataset including four labels: person (PER), miscellaneous (MISC), location (LOC), and organization (ORG). The specifics of the input text and the corresponding ground truth are provided in the second row of Table 1.
During reformulation, a single unstructured text containing all label-mention pairs is split into several sequences. Each new sequence’s output includes the count of mentions for a specified label (denoted as “entity type”), followed by the $n^{th}$ mention of that label (denoted as “¡mention n¿”). Note that the count of mentions and their respective indices are represented using corresponding digit tokens from the LLM’s vocabulary. Specifically, if there are no mentions, the model is trained to immediately predict the “¡eos¿” token, bypassing the need to predict mentions.
Therefore, in this example, one original training data is transformed into five new training data entries. These include two for predicting “LOC” (with 2 mentions), one for predicting “MISC” (with 1 mention), one for predicting “PER” (with 1 mention), and one for predicting “ORG” (with 0 mentions, directly predicting “¡eos¿”). Moreover, the number of mentions for each label and the text corresponding to each mention index can be easily obtained from the original ground truth, meaning that the number of new examples depends on the ground truth of that particular example.
With the newly reformulated training examples, we then apply the standard instruction tuning procedure. The model takes a sequence of text $t_{1},t_{2},\ldots,t_{T}$ consisting of input unstructured text and output structured label-mention pair. The optimization objective is cross-entropy loss $\mathcal{L}$ which can be defined as follows:
$\mathcal{L}=-\frac{1}{T}\sum_{i=1}^{T}\log P\left(t_{i}\mid t_{1},t_{2},\ldots,t_{i-1}\right)$ (1)
where $P\left(t_{i}\mid t_{1},t_{2},\ldots,t_{i-1}\right)$ represents the probability of $i^{th}$ token $t_{i}$ given the sequence of preceding tokens $t_{1},t_{2},\ldots,t_{i-1}$, as predicted by the model. Note that loss calculation begins from the number of mention tokens (i.e., texts enclosed by dashed-line frames). Theoretically, loss from text spans such as “¡mention n¿” could be ignored during this calculation, since they simply prompt the mention’s order, which does not necessarily need to be generated by the model. However, our ablation studies show that ignoring these texts has negligible impact on model performance, a point further discussed in Section 4.3. Therefore, we adhere to the standard instruction tuning procedure.
This reformulation allows the model to focus one label-mention pair at a time, shortening the generated length per sequence. More details are shown in Appendix B.
3.2 Inference of Label-Mention Pairs
Given a trained LLM, we propose a two-step inference approach: firstly, to predict the number of mentions for a specific label based on the prompt; and secondly, given the label and provided index to precisely identify the corresponding mention.
[…]
In practice, if there are sufficient GPU resources, the inference for the number of mentions for each label, as well as the subsequent inference for the mention text spans, can be allocating on separate GPUs. If GPU resources are limited, the inference can also be deployed on a single GPU using batch inference, facilitating parallel decoding. Using Figure 2 as an example, in Step 1, the batch size is four, as there are four labels in the dataset. In Step 2, the batch size is five, reflecting the five label-mention pairs determined in Step 1 (i.e., 1 in “PER”, 2 in “MISC”, 2 in “LOC”). This parallel decoding strategy is effective in reducing inference latency, especially in scenarios where inputs are received in a streaming manner.
3.3 Removal of Duplicate Mentions
Unlike autoregressive decoding, where subsequent label-mention pairs can attend preceding ones, PaDeLLM-NER generates each label-mention pair independently. This inference strategy means that the model might generate mentions erroneously repeated in multiple labels. As exemplified in Figure 2, the model correctly predicts the first mention of “LOC” as “Italy”, but it also incorrectly predicts the second mention of “MISC” as “Italy”.
To address the issue of duplicate mentions, we suggest employing prediction probability to remove repeated mentions. Specifically, we calculate the prediction probability for each instance of the mention. This is done using the formula: $P=\prod_{i=b}^{e}P(t_{i}|t_{1},t_{2},\ldots,t_{i-1})$ where $b$ represents the starting token index of the mention text, and $e$ denotes the ending token index. Then, for a mention that appears in multiple labels, the mention instance with the highest probability will be preserved. As illustrated in Figure2, “Italy” is categorized as “MISC” with only a 0.61 probability, which is lower than that for “LOC”, resulting in its removal. In practice, the probability of each token can be calculated concurrently with token generation. Consequently, this method enables an efficient and accurate identification of duplicate mentions without incurring additional costs. The effectiveness of this de-duplication approach is further explored in Section 4.3.
4 Experiments
In this section, we showcase the effectiveness of PaDeLLM-NER in terms of prediction quality and inference acceleration through experiments.
4.1 Setup
Datasets
We evaluate our method on English and Chinese NER datasets. English datasets include the general domain flat NER CoNLL2003 Tjong Kim Sang and De Meulder (2003), the nested NER ACE2005 Kirkpatrick (2010), and the biomedical nested NER GENIA Ohta et al. (2002). Chinese datasets include four commonly used general domain flat NER benchmarks Resume Zhang and Yang (2018), Weibo Peng and Dredze (2015), MSRA Levow (2006) and Ontonotes 4.0 Pradhan et al. (2013) and two vertical industrial domain flat NER datasets YouKu Jie et al. (2019) and Ecommerce Ding et al. (2019). The statistics of all datasets are shown in Appendix A.
Training setup
We employ pre-trained version of Llama2-7b Touvron et al. (2023b)***https://huggingface.co/meta-llama/Llama-2-7b and Baichuan2-7b Yang et al. (2023a)***https://huggingface.co/baichuan-inc/Baichuan2-7B-Base as base models for English and Chinese study respectively. Additional implementation details are in Appendix C.
Inference setup
For all generative models, we use greedy search with a beam size of $1$, a maximum of $512$ new tokens, and a temperature of $1.0$. As described in Section 3.2, for PaDeLLM-NER, we adopt two inference settings: (1) each example is inferred on multiple GPUs to implement parallel decoding (i.e., each sequence is assigned on one GPU), termed as PaDeLLMMulti; and (2) each example is inferred on a single GPU, employing batch decoding for parallel decoding, termed as PaDeLLMBatch. Note that for PaDeLLMMulti, we sequentially predict each sequence of one example to simulate parallel decoding on multiple GPUs.
Baselines
As the primary focus of this work is on reducing inference latency in NER tasks using LLMs, we compare our method, PaDeLLM-NER, with traditional autoregressive approaches. As mentioned in Section 1, the main points of comparison are autoregressive structured output formats used in Paolini et al. (2020); Das et al. (2023) and Lu et al. (2022, 2023a); Wang et al. (2023c), referred to respectively as AutoRegAug and AutoRegStruct, as these are the approaches very close to our system. We reimplemented these methods for both English and Chinese datasets, utilizing the same pre-trained LLMs as in PaDeLLM-NER. More details on the reimplementation are provided in Appendix C. Besides, we compare our approach with other recent state-of-the-art methods, including BINDER Zhang et al. (2022a), Gollie Sainz et al. (2023), and DeepStruct Wang et al. (2022a) for English benchmarks, as well as W2NER Li et al. (2022), NEZHA-BC Zhang et al. (2022b), and SSCNN Zhang and Lu (2023) for Chinese benchmarks, to show PaDeLLM-NER’s efficacy in prediction quality.
Evaluation
Our evaluation encompasses two dimensions: prediction quality and acceleration of NER inference. For assessing prediction quality, in line with Lu et al. (2022); Wang et al. (2023c), we employ the micro F-score.
Regarding inference acceleration, as per Ning et al. (2023), we evaluate using latency (in milliseconds). We record the latency using the following method: start = time.time(); model.generate(); latency = time.time() - start. In PaDeLLM-NER, we add the latency of mention counting and label-mention pair generation as the latency of each sequence. The final latency for the example is determined by the highest latency across sequences, as the user can only obtain the result of an example when the slowest sequence is generated. We conduct experiments three times and use the average result to alleviate the effect of randomness. We also report the average sequence length (tokenized) to clearly demonstrate the extent of sequence length reduction in Appendix D. Evaluations of all models were performed on the same NVIDIA A100 GPU.
4.2 Main Results
English Dataset Chinese Dataset
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Mean
AutoRegAug 992.70 944.90 1,515.35 1,276.32 812.78 1,009.68 982.39 579.99 845.42 995.50
AutoRegStruct 753.36 1,293.87 1,266.31 1,630.62 609.34 783.28 1,462.56 598.59 738.20 1,015.12
Ours
PaDeLLMMulti 229.74 255.53 316.90 159.57 143.47 171.67 238.27 203.63 293.40 223.57
[…]
Table 2: Comparison of inference latency (in milliseconds) between PaDeLLM-NER and baseline methods. Underscored font is the second-best method, while a bold font is the best method, also applied to subsequent tables.
(a) ARAug vs. PDLMMulti
(b) ARStruct vs. PDLMMulti
(c) ARAug vs. PDLMBatch
(d) ARStruct vs. PDLMBatch
Figure 3: Speedup of PaDeLLM-NER compared to Autoregressive methods.
English Dataset Chinese Dataset
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Mean
[…]
Ours
[…]
Table 3: Comparison of prediction quality between PaDeLLM-NER and baseline methods.
Latency evaluation
We investigate how PaDeLLM-NER reduces the end-to-end latency compared to baseline methods. Table 2 presents the average latency for each method across all datasets. First, it’s clear that both PaDeLLMMulti and PaDeLLMBatch significantly reduce inference latency when compared to baseline methods, as highlighted by the substantial reduction in mean latency. For example, the mean latency reduction achieved between PaDeLLMMulti and AutoRegStruct stands at an impressive $791.55$ ms, underscoring the significant improvement.
To more intuitively quantify the latency reduction of PaDeLLM-NER, we break down its speedup across different datasets in comparison to baseline methods in Figure 3. The speedup is computed by dividing the latency of baselines by the latency of PaDeLLM-NER. We can observe that PaDeLLM-NER consistently show a speedup over baseline methods across all datasets. The highest speedup is observed in the Weibo dataset when comparing AutoRegStruct vs. PaDeLLMMulti, with a speedup of $10.22$x. When we narrow our focus to the comparison between PaDeLLMBatch and the baseline methods, considering these methods utilize a single GPU for inference, we can still observe substantial speedup ranging from $1.76$x to $4.73$x. The speedup factor varies across different datasets, suggesting that the efficiency gains of PaDeLLM-NER may be influenced by the characteristics of each dataset. Interestingly, we can observe that the PaDeLLMBatch is slower than PaDeLLMMulti ($378.40$ ms vs. $223.57$ ms), more analysis about this is shown in Section 5.
Overall, the Table 2 and Figure 3 suggest that PaDeLLM-NER significantly reduces latency compared to autoregressive methods, though the extent of this reduction varies by dataset and the specific baseline method it’s compared to.
Prediction quality evaluation
Table 3 shows the micro F-score of PaDeLLM-NER and other baseline methods. It’s noteworthy that the micro F-scores for PaDeLLMMulti and PaDeLLMBatch are identical. Overall, PaDeLLM-NER outperforms baselines with the highest mean F-score of $84.79$ and emerging as the best-performing method in 4 out of the 9 datasets. It excels particularly in the Weibo, Youku and ACE2005 datasets and is competitive in others. We believe that the improved prediction could be attributed to shorter sequences, which lessen the challenge of long-range dependencies and enhance prediction quality. More analysis is discussed in Appendix E.
In summary, the results presented in Tables 2 and 3, demonstrate that our approach not only maintains superior prediction quality but also significantly reduces inference latency.
Comparison to state-of-the-art methods
In Tables 4 and 5, we compare the micro F-scores of PaDeLLM-NER with other recent state-of-the-art methods. Our findings indicate that PaDeLLM-NER performs comparably to these methods across most datasets except for Weibo. Notably, in the Youku dataset, PaDeLLM-NER outperforms the previously best-performing method by an improvement of 1.81%. Note that the primary focus of our work is on accelerating prediction latency using LLMs compared to baseline autoregressive methods, rather than achieving state-of-the-art results.
State-of-the-art Method CoNLL03 ACE05 GENIA
BINDER Zhang et al. (2022a) 93.33 89.50 80.50
Gollie Sainz et al. (2023) 93.10 89.60 -
DeepStruct Wang et al. (2022a) 93.00 86.90 80.80
Our Method
PaDeLLM-NER 92.52 85.02 77.66
Table 4: Comparing PaDeLLM-NER Performance with recent state-of-the-art methods on English datasets.
State-of-the-art Method Weibo MSRA Onto4 Resume Youku Ecom
NEZHA-BC Zhang et al. (2022b) - - - - - 82.98
SSCNN Zhang and Lu (2023) 71.81 - 82.99 96.40 86.10 81.80
W2NER Li et al. (2022) 72.32 96.10 83.08 96.65 - -
Our Method
PaDeLLM-NER 67.36 95.03* 80.81 94.98 87.91 81.85
Table 5: Comparing PaDeLLM-NER Performance with recent state-of-the-art methods on Chinese datasets. “*” indicates that results are not directly comparable.
4.3 Ablation study
Variant CoNLL03 ACE05 GENIA Mean
PaDeLLM-NER 92.52 85.02 77.66 85.06
+ Loss ignoring 92.01 85.18 73.47 83.55
[…]
+ De-duplicationReverse 92.38 84.44 77.38 84.73
Table 6: Ablations on ignoring loss and de-duplication.
In this section, we set out to investigate the effects of the different aspects of PaDeLLM-NER.
Ignoring text spans in loss
As discussed in Section 3.1, during training, it is permissible to overlook the loss of text span “¡mention n¿”, as the model does not need to generate this specific text, which is appended during inference. However, as shown in Table 6 illustrate, omitting these texts has minimal impact on prediction quality.
One possible explanation is that during training, the more significant challenge for LLMs lies in predicting the appropriate mention texts, rather than their format. As the model can readily learns to correctly position the format “¡mention n¿”, this aspect contributes minimally to the loss computation in training. In this case, computing the loss for all text is almost equivalent to ’neglecting’ the computation of loss for “¡mention n¿”.
De-duplication
To demonstrate the effectiveness of the de-duplication technique, we established two configurations as detailed in Table 6. The -De-duplication denotes the pipeline operating without the de-duplication technique; +De-duplicationReverse indicates the pipeline that removes mentions with the highest probability, opposite to the original de-duplication technique.
Theoretically, PaDeLLM-NER should be the top-performing method, as its de-duplication eliminates noisy mentions, enhancing precision. Following closely is the -De-duplication, allows duplicate mentions to persist. +De-duplicationReverse ranks lowest since it removes correct mentions and retains incorrect ones, lowering recall and precision simultaneously. As shown in Table 6, the results consistently align with our expectations, thereby verifying the effectiveness of the de-duplication process. Moreover, the difference among these variants is subtle, which can be attributed to the rare cases where duplicate mentions exist. This further highlights the robustness of proposed method.
5 Speedup Analysis
One concern noted is that batch inference does not speed up as much as inference distributed across multiple GPUs. This observation is consistent with our expectations and supported by Chen et al. (2023c) who found that batch inference in LLMs tends to be slower than single sequence inference under identical conditions, likely due to limitations in GPU memory bandwidth Cai et al. (2024).
Transitioning from these performance considerations, it’s noteworthy that PaDeLLM-NER is self-contained and can be seamlessly integrated with various generative architectures, including well-established decoder-only models Raffel et al. (2020a); Muennighoff et al. (2022); Touvron et al. (2023a, b); Bai et al. (2023); Yang et al. (2023a) and recent innovations like RWKV Peng et al. (2023), as well as multi-modal LLMs Feng et al. (2023); Liu et al. (2023b, a) for tasks like Key Information Extraction tasks Huang et al. (2019), all without needing architectural changes or additional data/modules. Also, it could be incorporated with off-the-shelf LLMs such as ChatGPT Achiam et al. (2023) and Claude-2***https://www.anthropic.com/news/claude-2 through prompt engineering without the need for further training, an aspect we plan to explore in future research.
6 Data Contamination Concerns
Since we are using LLMs as our foundational models, trained on extensive datasets from various online sources Touvron et al. (2023b); Yang et al. (2023a), there is a chance that the models may have encountered parts of our evaluation sets during their pre-training phase, albeit unintentionally. This could potentially affect our experimental results. However, the primary focus of our experiments is the comparison of our proposed method with baseline methods. Given that these methods employ the same LLM as the base model, data contamination is unlikely to significantly impact the results.
7 Conclusion
In this work, we introduce Parallel Decoding in LLM for NER (PaDeLLM-NER), a parallel decoding framework in LLMs for efficient NER. To achieve this, we recast autoregressive prediction of all label-mention pairs of traditional NER tasks into a two-step prediction: (1) predicting the number of mentions for a specific label, and (2) identifying the $n^{th}$ mention for that label. This recast allows the model to parallel decode all label-mention pairs in batches. Extensive experimental results show that the proposed method can dramatically reduces inference time, achieving inference speedup ranging from 1.76 to 10.22 times, without compromising prediction quality. Lastly, extensive ablation studies are performed to clarify the design choices of PaDeLLM-NER.
8 Limitations
One clear disadvantage of PaDeLLM-NER is the multiplication of training examples from one to $m*n$, where $m$ is the label count and $n$ the mention count. Despite this, given that low latency is a major bottleneck in LLMs, trading longer training for lower latency is justifiable. Also, given the impressive generalization ability of LLMs, we believe that this method can be smoothly adapted to few-shot scenarios requiring less computation resources, which will be explored in future work.
Additionally, accurately counting the number of mentions remains a challenge for LLMs as discussed in Appendix E. This issue could be alleviated by implementing a specialized counting model dedicated to this task Liu and Low (2023).
Finally, there are several instances of re-computation within the pipeline that can be optimized. Specifically, input texts are encoded multiple times throughout the process. During batch decoding, certain sequences may encounter the “¡eos¿” token earlier, but due to the nature of batch inference, these sequences continue to predict. We plan to improve this in the future by implementing enhancements like KV cache reuse and batch inference with an early quit mechanism, among other strategies.
Appendix A Dataset Statistics
We evaluate our framework on 3 English and 6 Chinese flat/nested NER datasets. In Table 9, we present the detailed statistics. Note that while the statistics of the development set are reported, our training process does not involve the development set.
For the MSRA dataset, we excluded four outlier instances from the test set due to their excessively high number of mentions, significantly deviating from typical examples. These outliers not only posed challenges for model inference but also risked distorting the evaluation metrics, potentially leading to an inaccurate assessment of the model’s performance on representative data.
Also, we perform label mapping to convert ground truth from special tokens to Chinese words following Lu et al. (2023a). Further details are provided in Table 10.
Dataset Sentence Mention
#All #Train #Dev #Test #All #Train #Dev #Test
CoNLL2003 20,744 14,041 3,250 3,453 35,089 23,499 5,942 5,648
ACE2005 9,210 7,194 969 1,047 30,634 24,441 3,200 2,993
GENIA 18,546 15,023 1,669 1,854 56,015 46,142 4,367 5,506
Weibo 1,890 1,350 270 270 2,701 1,894 389 418
MSRA* 50,725 44,364 - 4,361 80,214 74,703 - 5,511
OntoNotes 4.0 24,371 15,724 4,301 4,346 28,006 13,372 6,950 7,684
Resume 4,759 3,819 463 477 16,565 13,438 1,497 1,630
Youku 10,002 8,001 1,000 1,001 15,905 12,754 1,581 1,570
Ecommerce 4,987 3,989 500 498 15,216 12,109 1,540 1,567
Table 9: Dataset Statistics. “#” denotes the amount. For MSRA, we remove four outlier examples in test set.
Dataset #Entity Entity
Weibo 8
{“PER.NAM(Specific Name)”:“名称特指”, “PER.NOM(Generic Name)”:“名称代称”, “GPE.NAM(Specific Geo-Political Entity)”:“行政区特指”, “GPE.NOM(Generic Geo-Political Entity)”:“行政区代称”, “LOC.NAM(Specific Location)”:“地点特指”, “LOC.NOM(Generic Location)”:“地点代称”, “ORG.NAM(Specific Organization)”:“组织特指”, “ORG.NOM(Generic Organization)”:“组织代称” }
MSRA 3
{“LOC”:“地点’, “PER”:“名称”, “ORG”:“组织”}
OntoNotes 4.0 4
{“GPE”:“地缘”, “LOC”:“地点”, “PER”:“名称”, “ORG”:“组织”}
Resume 8
{“NAME”:“名称”, “CONT(Nationality)”:“国籍”, “RACE”:“民族”, “TITLE”:“职位”, “EDU”:“学历”, “ORG”:“公司”, “PRO(Profession)”:“专业”, “LOC(Place of Birth)”:“籍贯”}
Youku 3
{“TELEVISION”:“电视剧”, “PER(Celebrity)”:“明星”, “MISC”:“其他”}
Ecommerce 2
{“HP(brand)”:“品牌”, “HC(commodity)”:“商品”}
Table 10: Entity tag of each dataset and the conversion from tag used in dataset to corresponding Chinese natural language. For some tags that are hard to understand, we provide their meaning in brackets. “#” denotes the amount of entity types.
Appendix B Reformulation Examples
Two compete reformulated examples are presented in Table 11 for English and Chinese, respectively.
Language Input Output
English text: But Fischler agreed to review his proposal after the EU ’s standing veterinary committee , mational animal health officials , questioned if such action was justified as there was only a slight risk to human health . entity type: PER ¡num¿ 1 ¡mention 1¿Fischler
Chinese 文本(text): 公报最后说,墨西哥政府认为,贩毒以及洗钱等与毒品有关的活动是威胁到国家主权和安全的一个全球性问题。(The communique concluded by stating that the Mexican government considers drug trafficking and related activities such as money laundering to be a global issue that threatens national sovereignty and security.) 指定NER标签(entity type): 地点(LOC) ¡数量¿(¡num¿) 1 ¡第1文段¿(¡mention 1¿) 墨西哥(Mexican)
Table 11: Reformulated examples for English and Chinese dataset, respectively. We provide translations to facilitate understanding. The examples come from CoNLL2003 and MSRA dataset.
Appendix C Implementation Details
We train our model on all datasets for $4$ epochs, using a batch size of $128$ and a learning rate of $51e-$ 5, with the AdamW optimizer Loshchilov and Hutter (2018) and a cosine scheduler Loshchilov and Hutter (2017). The maximum input and output sequence lengths are set to 2048 and 512, respectively. Training is conducted on 8 NVIDIA A100 GPUs. This configuration is applied across all PaDeLLM-NER models, as well as two baseline models: AutoRegAug and AutoRegStruct.
Appendix D Sequence Length Reduction
Results of average sequence length produced by different approaches are presented in Table 8. Most notably, PaDeLLM-NER generates much shorter sequences than the other models across all datasets. The lengths range from 6.54 on CoNLL20023 to 10.05 on GENIA for English datasets, and from 2.19 on Weibo to 4.87 on Resume for Chinese datasets. The mean length for PaDeLLM-NER is 4.86, which is significantly lower than the means of the other approaches: 35.54 for AutoRegAug and 36.48 for AutoRegStruct.
In summary, the result shows that PaDeLLM-NER produces much shorter generated sequences compared to the other methods, which is around 13.19% to 13.67% of the original length, respectively, indicating higher efficiency in its inference.
Appendix E Error analysis
PaDeLLM-NER error analysis
For our error analysis, we utilize the ACE2005 dataset. We sample and manually examine 50 erroneous examples for analysis. We seek to identify the root causes of errors, which we have categorized into three types: (1) incorrect mention count, referred to as Count Mismatch; (2) inaccuracies in the mention corresponding to a specific index, termed Index Inaccuracy; and (3) errors in the ground truth data, known as Ground Truth Errors.
The distribution of each error type is illustrated in Figure 4. It is important to note that a significant portion of the errors stem from inaccuracies in mention counts (i.e., Count Mismatch, about 56.8%), underscoring the necessity for enhancements in the model’s counting capabilities. Accurate mention counts are pivotal for the quality of predictions. Overestimating the mention count often leads the model to either repeat the last entity or, more problematically, fabricate an entity, thereby escalating the rate of false positives. Conversely, underestimating the mention count results in the model’s inability to identify some entities, thus increasing the incidence of false negatives. Following closely is the Index Inaccuracy error, indicating that the model sometimes struggles to accurately pinpoint the correct mention for a given index, further emphasizing areas for improvement.
Interestingly, our analysis reveal that a significant portion of the model’s predictions, specifically 19.3%, are actually correct, challenging the accuracy of the ground truth data. This observation suggests the presence of inaccuracies within the ground truth, contributing to an elevated rate of false positives. Prior research, as noted in studies by Min et al. (2022); Wang et al. (2023a); Zhou et al. (2023), has demonstrated that LLMs predominantly acquire their knowledge during the pre-training phase. These models develop certain “core beliefs” that tend to align more closely with human judgment. In this context, it appears that the models possess an inherent capability to rectify errors in the ground truth data, demonstrating their potential to improve data accuracy beyond initial human annotation.
Appendix F Model Scaling Up
As we increase the model size to 13B, Table 7 presents a mix of results. In datasets like CoNLL2003 and GENIA, the model shows a significant improvement in predictions. In contrast, the results on ACE2005 are slightly worse. Note that the improvement in GENIA is substantial, at approximately 1.18%. Based on these findings, it seems reasonable to suggest that continuously scaling up the model size has the potential to maintain the performance that is at least on par, or even superior, especially in specific industrial domains. However, this hypothesis warrants further investigation, involving more families of models Raffel et al. (2020a); Muennighoff et al. (2022); Touvron et al. (2023a, b); Bai et al. (2023); Yang et al. (2023a) and a broader range of datasets. We leave this exploration for future work.